# Unifying AI Tutor Evaluation: An Evaluation Taxonomy for Pedagogical Ability Assessment of LLM-Powered AI Tutors

### v1

@@ Abstract
In this paper, we investigate whether current state-of-the-art large language models (LLMs) are effective as AI tutors and whether they demonstrate pedagogical abilities necessary for good AI tutoring in educational dialogues. Previous efforts towards evaluation have been limited to subjective protocols and benchmarks. To bridge this gap, we propose a unified evaluation taxonomy with eight pedagogical dimensions based on key learning sciences principles, which is designed to assess the pedagogical value of LLM-powered AI tutor responses grounded in student mistakes or confusion in the mathematical domain. We release MRBench – a new evaluation benchmark containing 192 conversations and 1,596 responses from seven state-of-the-art LLM-based and human tutors, providing gold annotations for eight pedagogical dimensions. We assess reliability of the popular Prometheus2 LLM as an evaluator and analyze each tutor’s pedagogical abilities, highlighting which LLMs are good tutors and which ones are more suitable as question-answering systems. We believe that the presented taxonomy, benchmark, and human-annotated labels will streamline the evaluation process and help track the progress in AI tutors’ development.
@@
Kaushal Kumar Maurya, KV Aditya Srivatsa, Kseniia Petukhova and Ekaterina Kochmar Mohamed bin Zayed University of Artificial Intelligence, Abu Dhabi, UAE {kaushal.maurya, vaibhav.kuchibhotla, kseniia.petukhova, ekaterina.kochmar}@mbzuai.ac.ae
@@ 1 Introduction
Human tutoring is a cornerstone of educational development, playing a crucial role in fostering societal growth by empowering learners. While one-on-one tutoring is highly effective Bloom (1984), its ubiquitous implementation is hindered by the limited availability of qualified tutors. [^1] https://unesdoc.unesco.org/ark:/48223/pf0000385723 The remarkable success of LLMs as conversational systems offers promising opportunities in education Wang et al. (2024b); Gan et al. (2023), driving the development of LLM-powered intelligent tutoring systems (ITS) (Pal Chowdhury et al., 2024; Liu et al., 2024)) and the deployment of LLMs as tutors using advanced prompting techniques Denny et al. (2024); Mollick and Mollick (2024). Such AI tutors serve various educational objectives Wollny et al. (2021), among which the task of students’ mistake and confusion remediation is one of the most popular, leading to active AI tutor development Macina et al. (2023); Wang et al. (2024a).
@@ 2.2 Pedagogically-oriented Evaluation
Most of the traditional evaluation methods from learning sciences are designed for the evaluation of human tutors and can not be easily applied to AI tutors due to the absence of a self-play report Tack and Piech (2022). A reliable avenue is to hire human experts to evaluate pedagogical performance Vasselli et al. (2023); Lee et al. (2023); Abdelghani et al. (2024). However, there is no agreed-upon protocol for conducting pedagogical human evaluations, and researchers consider various pedagogical dimensions and their associated definitions Wollny et al. (2021); Tack et al. (2023); Denny et al. (2024). The most commonly used evaluation framework involves human raters comparing the responses of two tutors in the context of the same dialogue snippet Tack and Piech (2022). These comparisons are based on the three dimensions defined by Demszky et al. (2021), but they do not fully capture pedagogical richness. Similar efforts by Macina et al. (2023); Wang et al. (2024a); Daheim et al. (2024) partially cover pedagogical aspects in the student mistake remediation task in the mathematical domain. A large-scale study by Jurenka et al. (2024) concluded that there is a need to develop well-recognized, unified evaluation metrics that enable comparisons across different models and track the progress of AI tutors. Our proposed taxonomy is a step toward this ambitious goal, and we believe this effort will streamline the evaluation of AI tutors and their pedagogical abilities.
@@ 3 Student Mistake Remediation Task
Formally, we define the conversation history between a tutor and a student as =$\mathcal{H}=\{(\mathcal{T}_{1},\mathcal{S}_{1}),(\mathcal{T}_{2},\mathcal{S}_{2}),\ldots,(\mathcal{T}_{t},\mathcal{S}_{t})\}$, where $\mathcal{T}_{i}$ represents the $i$-th response from the tutor and $\mathcal{S}_{i}$ represents the $i$-th response from the student. Let $\mathcal{S}_{t}$ denote the student’s last utterance containing a mistake or confusion. Then the objective of the tutor is to provide the most appropriate response $\mathcal{T}_{t+1}$ to address this mistake or confusion. The evaluation taxonomy detailed in Section 4 assesses the appropriateness of the $\mathcal{T}_{t+1}$ response across eight key pedagogical dimensions.
@@
Mistake identification Has the tutor identified a mistake in a student’s response? Yes
@@
Coherence Is the tutor’s response logically consistent with the student’s previous response? Yes
@@ 4.1 Evaluation Taxonomy Dimensions
Overall, our schema covers all the relevant aspects of a good tutor response proposed in previous work Tack and Piech (2022); Macina et al. (2023); Wang et al. (2024a); Daheim et al. (2024) while also being supported by the learning sciences principles. Although there are inherent inter-dependencies among the proposed dimensions of the taxonomy (e.g., revealing the answer is less likely to be actionable, and vice versa), we explicitly instructed all annotators to treat each dimension as independent and orthogonal to minimize confounding factors and potential biases during the annotation process.
@@ 4.2 Evaluation Taxonomy Validation
To evaluate the efficacy and pertinence of the proposed evaluation taxonomy, we conducted a series of validation experiments aimed at addressing the critical question: Are the proposed dimensions sufficient, and is there redundancy among them? The validation process comprised the validation pilot study involving four annotators who iteratively reviewed the annotation scheme and guidelines and independently labeled eight randomly sampled dialogues (four from each of the two datasets introduced in Section 5.1). All four annotators possess at least a post-graduate degree and are proficient in English. The annotation team consisted of two male and two female annotators. To control the annotation workflow and ensure quality, we opted not to use public annotation outsourcing platforms such as Prolific or MTurk. This allowed us to implement rigorous training protocols and a robust validation mechanism for the annotations. We provided all annotators with comprehensive training, including an interactive training document (see Section C for more detail) and oral instructions. Following this, we conducted a pilot test to evaluate the annotators’ understanding and annotation quality before rolling out the large-scale human evaluation detailed in Section 5.1. This multi-step process ensured that the annotations adhered to our quality standards. In this validation pilot study, all four annotators assessed the responses across the eight dimensions of evaluation taxonomy. To measure inter-annotator agreement, we computed Fleiss’ kappa value, which for this annotation equals 0.65, indicating substantial agreement. None of the annotators identified any additional or redundant dimensions necessary for student mistake remediation.
@@ 5 Pedagogical Ability Assessment Settings
In this section, we provide details on the benchmark data preparation and statistics, deployed LLMs as AI tutor, the human annotation process, and the LLM-based evaluation.
@@ 5.1 Benchmark Preparation
The Bridge dataset Wang et al. (2024a) comprises partial dialogue interactions between real human tutors and students at the elementary level, featuring two distinct human tutor responses (novice and expert). The dialogue context is typically short (few turns) and predominantly focused on fundamental mathematical concepts, including operations such as multiplication, addition, and so on. The original dataset consists of a total of 700 dialogues; we filtered 60 high-quality instances for MRBench. [^2] Many examples in the dataset are grounded in visual contexts, which we have not incorporated into this benchmark.
The dialogues in the MathDial dataset Macina et al. (2023) consist of complete multi-turn conversations between a real human tutor and an LLM acting as a student, where the tutor aims to remediate the student’s mistakes. Specifically, these conversations are grounded in middle school-level mathematical reasoning questions. To match the format of Bridge (partial conversations with the last student’s utterance exhibiting a mistake or confusion), we prepared the dataset by terminating a conversation where the student makes a mistake and considering the next tutor response as the expert tutor response (there are no novice responses in this dataset). To further ensure the reliability of our benchmark, we manually inspected to retain only high-quality examples, resulting in 132 instances for MRBench.
Next, for the 192 instances in MRBench (60 from Bridge and 132 from MathDial), we generated appropriate subsequent responses based on the conversation history and the last utterance, which contained confusion or mistakes, using seven state-of-the-art LLMs. These models were prompted to act as expert tutors (see Figure 2 for the exact prompt template). We consider state-of-the-art LLMs of various sizes and capabilities, including: GPT-4 Achiam et al. (2023), Gemini Reid et al. (2024), Sonnet Anthropic (2024), Mistral Jiang et al. (2023), Llama-3.1-8B and Llama-3.1-405B Dubey et al. (2024), and Phi3 Abdin et al. (2024).
@@ 5.3 LLM-based Annotation
Due to the growing interest in utilizing LLMs as critics or evaluators Jurenka et al. (2024); Chang et al. (2024), we also used a separate LLM as an evaluator. Specifically, we used Prometheus2 Kim et al. (2024) because: (i) it was specifically trained as an evaluator using reinforcement learning with human feedback (RLHF), (ii) it has a high correlation with human annotations and GPT-4, and (iii) it does not belong to any of the LLM families considered as AI tutors in our framework.
@@ 6 Key Findings
This section summarises the key findings of our study on the pedagogical abilities of LLMs as AI tutors, based on human and LLM-based evaluation of MRBench, and the correlation between them. We consider human-based evaluations as gold standard. Table 3 shows DAMR scores for each LLM across all eight dimensions.
@@
To account for diversity, we also included two lightweight LLMs (with fewer parameters) as tutors, namely Llama-3.1-8B and Phi3. Phi3 is the worst-performing LLM model in this context, with the lowest score for coherence, suggesting that the responses from Phi3 are often irrelevant to the conversation context, as well as overall low scores in other dimensions. This underscores the model’s inadequate capacity for contextual understanding and semantic alignment. In the few cases where Phi3 demonstrates some competency, it frequently reveals the answer, reflecting more of a question-answer system rather than a pedagogical tutor behavior. Moreover, its outputs tend to be robotic, template-based and lack the nuance expected in human responses. Conversely, despite having fewer parameters, Llama-3.1-8B demonstrates reasonable performance, albeit still below that of larger LLMs. Specifically, its responses are coherent, strategically avoid immediate answer revelation, robustly identify and rectify mistakes, and exhibit human-like behavior, as evidenced by the DAMR scores.
@@
We also investigated the pedagogical value of human responses for both Novice and Expert. It can be observed that Novice responses do not have a high score for guidance and are poor in terms of actionability (DAMR score of 1.67). Furthermore, the responses are generally short and ambiguous, such as "this is a good try," which does not reflect human-likeness and also leads to lower scores for mistake identification and location. At the same time, they often do not reveal the answer. In contrast, Expert human responses are more logical and highly actionable for the next steps. However, in a few cases, their responses are more action-oriented even if they do not provide factually correct guidance, leading to lower scores for guidance compared to actionability. This leads to a question: Can a tutor achieve a higher DAMR score for actionability while receiving a lower score for providing guidance? This is possible since we consider only factually correct guidance as useful (see Table 4). At the same time, even incorrect or incomplete guidance can lead to certain actions on the part of the student and can foster their curiosity, thus providing them with learning opportunities. For example, a response like "24 x 10 = ?" does not provide guidance, yet it is actionable. This further demonstrates the need to treat the dimensions as independent. In terms of the other qualities of the Expert responses, they do not normally reveal the answer and tend to include scaffolding; however, there are a small number of instances where they failed to identify the mistake or its location. Overall, we conclude that human responses from Expert are significantly better than Novice.
@@
Our findings on the Tutor Tone align with those of Wang et al. (2024a) – in task-oriented conversations, AI tutors tend to be more Neutral than Encouraging. When we combine these two labels into "Non-offensive", the DAMR score reaches 100% as we observe no offensive responses from any LLMs or humans. We observe high scores for most of the LLMs on human-likeness, which demonstrates their capability to generate human-like output with minimal or no grammatical and fluency mistakes, showing the timely nature of our study, which focuses more on in-depth semantic and pedagogical aspects of tutor responses rather than only superficial attributes like grammaticality and fluency.
@@
As discussed in Section 5.1, the conversational contexts in the Bridge dataset are typically very short (see Table 6) and the dialogues are grounded in elementary math operations mistakes, so most models are able to identify the mistakes and their locations. However, they struggle to provide appropriate guidance without revealing the answer because the mistakes are generally related to quite basic operations like addition or multiplication, often in a one-step type of mathematical problems. Still, models like GPT-4 and Llama-3.1-405B are able to offer some reasonable guidance. In contrast, for MathDial, the contexts are longer, the mistakes are grounded in reasoning, and the responses are more structured. Yet, many LLMs do not meet the expectations for each dimension of the taxonomy, as discussed earlier. DAMR scores for Bridge and Mathdial are shown in Appendix Table 7 and 8, respectively. Combining both types of data in MRBench makes it both challenging and comprehensive.
@@
In summary, all LLMs and even human tutors lack some pedagogical abilities required for effective tutoring. While Llama-3.1-405B is the most effective, followed by Mistral and other state-of-the-art models, GPT-4 reveals the answer too quickly. Gemini is less coherent and accurate, and Sonnet focuses on human-likeness and encouraging tone but is less effective in other dimensions. Phi3 is the worst-performing model according to our analysis, as it fails to understand the context, while Llama-3.1-8B, despite being smaller, performs reasonably well. Human responses are also not perfect – Novice responses are ambiguous and short, whereas Expert responses are more focused on actionability and less on other dimensions. Overall, the proposed taxonomy precisely categorizes performance across 8 dimensions, reflecting the current state-of-the-art in AI tutors. Our study demonstrates that there is a huge opportunity for improvement in the pedagogical abilities of AI tutors.
@@
We also performed the annotation with the popular critic LLM Prometheus2. The correlation scores with human annotations are presented in Appendix Table 5. It can be observed that most of the correlation scores are negative (except for the human-likeness dimension), which indicates that the annotations from the LLM are unreliable for the challenging pedagogical dimensions. There may be several reasons for that: (i) The Prometheus2 model is not trained on our taxonomy dimensions, except for general human-likeness, where it shows slightly better correlation with positive scores. However, the score for human-likeness is still low, as Prometheus2 requires gold-standard responses, which are not unique and were unavailable in our case. (ii) We believe, LLMs have limited understanding of rich pedagogical concepts, as they were not specifically trained on pedagogically rich datasets.

### v2

@@ Abstract
In this paper, we investigate whether current state-of-the-art large language models (LLMs) are effective as AI tutors and whether they demonstrate pedagogical abilities necessary for good AI tutoring in educational dialogues. Previous efforts towards evaluation have been limited to subjective protocols and benchmarks. To bridge this gap, we propose a unified evaluation taxonomy with eight pedagogical dimensions based on key learning sciences principles, which is designed to assess the pedagogical value of LLM-powered AI tutor responses grounded in student mistakes or confusions in the mathematical domain. We release MRBench – a new evaluation benchmark containing 192 conversations and 1,596 responses from seven state-of-the-art LLM-based and human tutors, providing gold annotations for eight pedagogical dimensions. We assess reliability of the popular Prometheus2 and Llama-3.1-8B LLMs as evaluators and analyze each tutor’s pedagogical abilities, highlighting which LLMs are good tutors and which ones are more suitable as question-answering systems. We believe that the presented taxonomy, benchmark, and human-annotated labels will streamline the evaluation process and help track the progress in AI tutors’ development.
@@
Kaushal Kumar Maurya, KV Aditya Srivatsa, Kseniia Petukhova and Ekaterina Kochmar Mohamed bin Zayed University of Artificial Intelligence {kaushal.maurya, vaibhav.kuchibhotla, kseniia.petukhova, ekaterina.kochmar}@mbzuai.ac.ae
@@ 1 Introduction
Human tutoring is a cornerstone of educational development, playing a crucial role in fostering societal growth by empowering learners. While one-on-one tutoring is highly effective Bloom (1984), its ubiquitous implementation is hindered by the limited availability of qualified tutors. [^1] https://unesdoc.unesco.org/ark:/48223/pf0000385723 The remarkable success of LLMs as conversational systems offers promising opportunities in education Wang et al. (2024b); Gan et al. (2023), driving the development of LLM-powered intelligent tutoring systems (ITS) (Pal Chowdhury et al., 2024; Liu et al., 2024) and the deployment of LLMs as tutors using advanced prompting techniques Denny et al. (2024); Mollick and Mollick (2024). Such AI tutors serve various educational objectives Wollny et al. (2021), among which the task of students’ mistake and confusion remediation is one of the most popular, leading to active AI tutor development Macina et al. (2023); Wang et al. (2024a).
@@ 2.2 Pedagogically-oriented Evaluation
Most of the traditional evaluation methods from learning sciences are designed for the evaluation of human tutors and can not be easily applied to AI tutors due to the absence of a self-reports Tack and Piech (2022). A reliable avenue is to hire human experts to evaluate pedagogical performance Vasselli et al. (2023); Lee et al. (2023); Abdelghani et al. (2024). However, there is no agreed-upon protocol for conducting pedagogical human evaluations, and researchers consider various pedagogical dimensions and their associated definitions Wollny et al. (2021); Tack et al. (2023); Borges et al. (2024); Denny et al. (2024). The most commonly used evaluation framework involves human raters comparing the responses of two tutors in the context of the same dialogue snippet Tack and Piech (2022). These comparisons are based on the three dimensions defined by Demszky et al. (2021), but they do not fully capture pedagogical richness. Similar efforts by Macina et al. (2023); Wang et al. (2024a); Daheim et al. (2024) partially cover pedagogical aspects in the student mistake remediation task in the mathematical domain. A large-scale study by Jurenka et al. (2024) concluded that there is a need to develop well-recognized, unified evaluation metrics that enable comparisons across different models and track the progress of AI tutors. Our proposed taxonomy is a step toward this ambitious goal, and we believe this effort will streamline the evaluation of AI tutors and their pedagogical abilities.
@@ 3 Student Mistake Remediation Task
Formally, let’s define the conversation history between a tutor and a student as =$\mathcal{H}=\{(\mathcal{T}_{1},\mathcal{S}_{1}),(\mathcal{T}_{2},\mathcal{S}_{2}),\ldots,(\mathcal{T}_{t},\mathcal{S}_{t})\},$ where $\mathcal{T}_{i}$ represents the $i$-th response from the tutor, and $\mathcal{S}_{i}$ represents the $i$-th response from the student. Let $\mathcal{S}_{k}$ denote the student’s most recent $k$ utterances, where $k\in[1,...,t]$, containing a mistake or confusion. Then the objective of the tutor is to provide the most appropriate response $\mathcal{T}_{t+1}$ to address this mistake or confusion. The evaluation taxonomy detailed in Section 4 assesses the appropriateness of the $\mathcal{T}_{t+1}$ response across eight key pedagogical dimensions.
@@
Mistake identification Has the tutor identified/recognized a mistake in a student’s response? Yes
@@
Coherence Is the tutor’s response logically consistent with the student’s previous responses? Yes
@@ 4.1 Evaluation Taxonomy Dimensions
Overall, our schema covers all the relevant aspects of a good tutor response proposed in previous work Tack and Piech (2022); Macina et al. (2023); Wang et al. (2024a); Daheim et al. (2024) while also being supported by the learning sciences principles. Although there are inherent inter-dependencies among the proposed dimensions of the taxonomy (e.g., a response that reveals the answer is less likely to be actionable, and vice versa), we explicitly instructed all annotators to treat each dimension as independent and orthogonal to minimize confounding factors and potential biases during the annotation process. Estimation of the relative importance among evaluation dimensions is beyond the scope of this study and is left for future work.
@@ 4.2 Evaluation Taxonomy Validation
To evaluate the efficacy and pertinence of the proposed evaluation taxonomy, we conducted a series of validation experiments aimed at addressing the critical questions: Are the proposed dimensions sufficient? and Are there redundancies among them? The annotation team consisted of two male and two female annotators, with all four annotators holding at least a post-graduate degree in Computer Science and being proficient in English. We note that for this study, we do not require annotators to have direct teaching experience, as understanding of the mathematical tasks at the middle school level and being able to judge the responses from the perspective of a potential user of such AI tutors (or a student), rather than specifically a teacher, is sufficient. To control the annotation workflow and ensure quality, we opted not to use public annotation outsourcing platforms such as Prolific or MTurk, which allowed us to implement rigorous training protocols and a robust validation mechanism for the annotations.
First, we provided all annotators with comprehensive training, including an interactive training document (see Section C for more details) and oral instructions. Following this, we conducted validation pilot study to evaluate the annotation quality and the annotators’ understanding of the instructions before rolling out the large-scale human evaluation detailed in Section 5.2. This multi-step process ensured that the annotations adhered to our quality standards. In this validation pilot study, all four annotators iteratively reviewed the annotation scheme and guidelines. Each annotator also independently labeled the same eight randomly sampled dialogues – four from each of the two datasets (Bridge and MathDial) – across the eight dimensions of the evaluation taxonomy. Given that each dialogue contained multiple responses from both LLMs and humans, and each response was annotated across eight evaluation dimensions, this resulted in a total of 544 annotations per annotator. To measure inter-annotator agreement, we computed Fleiss’ kappa value, which for this annotation experiment equals 0.65, indicating substantial agreement. None of the annotators identified any additional or redundant dimensions necessary for student mistake remediation.
@@ 5 Pedagogical Ability Assessment Settings
In this section, we provide details on the benchmark data preparation and statistics, LLMs deployed as AI tutors, the human annotation process, and the LLM-based evaluation.
@@ 5.1 Benchmark Preparation
The Bridge dataset Wang et al. (2024a) comprises partial dialogue interactions between real human tutors and students at the elementary level, featuring two distinct human tutor responses (novice and expert). The dialogue context is typically short (few turns) and predominantly focused on fundamental mathematical concepts, including operations such as multiplication, addition, and so on. The original dataset consists of a total of 700 dialogues; we filtered 60 high-quality instances for MRBench. [^2] Many examples in the dataset are grounded in visual contexts, which we have not incorporated into this benchmark. Among the various criteria for selecting high-quality dialogues, the key one was that the student’s last utterance (or last few utterances) should exhibit an error or confusion.
The dialogues in the MathDial dataset Macina et al. (2023) consist of complete multi-turn conversations between a real human tutor and an LLM acting as a student, where the tutor aims to remediate the student’s mistakes. Specifically, these conversations are grounded in middle school-level mathematical reasoning questions. To match the format of Bridge (partial conversations with the last few student’s utterances exhibiting a mistake or confusion), we prepared the dataset by terminating a conversation where the student makes a mistake and considering the next tutor response as the expert tutor response (there are no associated novice responses in this dataset). To further ensure the reliability of our benchmark, we manually inspected the data in order to retain only high-quality examples, which resulted in 132 instances for MRBench.
Next, for the 192 instances in MRBench (60 from Bridge and 132 from MathDial), we generated appropriate subsequent responses based on the conversation history and the last utterance, which contained confusions or mistakes, using seven state-of-the-art LLMs. These models were prompted to act as expert tutors (see Figure 2 for the exact prompt template). We consider state-of-the-art LLMs of various sizes and capabilities, including: GPT-4 Achiam et al. (2023), Gemini Reid et al. (2024), Sonnet Anthropic (2024), Mistral Jiang et al. (2023), Llama-3.1-8B and Llama-3.1-405B Dubey et al. (2024), and Phi3 Abdin et al. (2024).
@@ 5.3 LLM-based Annotation
Due to the growing interest in utilizing LLMs as critics or evaluators Jurenka et al. (2024); Chang et al. (2024), we also used two LLMs as evaluators:
  • We used Prometheus2 Kim et al. (2024) because: (i) it was specifically trained as an evaluator using reinforcement learning with human feedback (RLHF), (ii) it has a high correlation with human annotations and GPT-4, and (iii) it does not belong to any of the LLM families considered as AI tutors in our framework.
  • In addition, we also used Llama-3.1-8B as a lightweight LLM to assess the reliability of smaller models that were not fine-tuned for evaluation objectives as a critic.
@@ 6 Key Findings
This section summarizes the key findings of our study on the pedagogical abilities of LLMs as AI tutors, based on human and LLM-based evaluation of MRBench, and the correlation between them. We consider human-based evaluations as gold standard. Table 3 shows DAMR scores for each LLM across all eight dimensions.
@@
To account for diversity, we also included two lightweight LLMs (with fewer parameters) as tutors, namely Llama-3.1-8B and Phi3. Phi3 is the worst-performing LLM model in this context, with the lowest score for coherence, suggesting that the responses from Phi3 are often irrelevant to the conversation context, as well as overall low scores in other dimensions. This underscores the model’s inadequate capacity for contextual understanding and semantic alignment in educational dialogues considered in this study. In the few cases where Phi3 demonstrates some competence, it frequently reveals the answer, reflecting more of a question-answer system than a pedagogical tutor behavior. Moreover, its outputs tend to be robotic, template-based and lack the nuance expected in human responses. In contrast, despite having fewer parameters, Llama-3.1-8B demonstrates reasonable performance, albeit still below that of larger LLMs. Specifically, its responses are coherent, strategically avoid immediate answer revelation, robustly identify and rectify mistakes, and exhibit human-like behavior, as evidenced by the DAMR scores.
@@
We also investigated the pedagogical value of human responses for both Novice and Expert. It can be observed that Novice responses do not have a high score for guidance and are poor in terms of actionability (DAMR score of 1.67). Furthermore, the responses are generally short and ambiguous, such as "this is a good try," which leads to lower scores for mistake identification and location. At the same time, they often do not reveal the answer. In contrast, Expert human responses are more logical and highly actionable for the next steps. However, in a few cases, their responses are action-oriented even though they do not provide factually correct guidance, leading to lower scores for guidance compared to actionability. This leads to a question: Can a tutor achieve a higher DAMR score for actionability while receiving a lower score for providing guidance? This is possible since we consider only factually correct guidance as useful (see Table 4). At the same time, even incorrect or incomplete guidance can lead to certain actions on the part of the student and can foster their curiosity, thus providing them with learning opportunities. For example, a response like "24 x 10 = ?" does not provide guidance, yet it is actionable. This further demonstrates the need to treat the dimensions as independent. In terms of the other qualities of the Expert responses, they do not normally reveal the answer and tend to include scaffolding; however, there are a small number of instances where they failed to identify the mistake or its location. Overall, we conclude that human responses from Expert are significantly better than Novice.
@@
Our findings on the Tutor Tone align with those of Wang et al. (2024a) – in task-oriented conversations, AI tutors tend to be more Neutral than Encouraging. When we combine these two labels into "Non-offensive", the DAMR score reaches 100% as we observe no offensive responses from any LLMs or humans. We observe high scores for most of the LLMs on human-likeness, which demonstrates their capability to generate human-like output with minimal or no grammatical and fluency mistakes, showing the timely nature of our study, which focuses more on in-depth semantic and pedagogical aspects of tutor responses rather than only on superficial attributes like grammaticality and fluency.
@@
As discussed in Section 5.1, the conversational contexts in the Bridge dataset are typically very short (see Table 7) and the dialogues are grounded in elementary math operations, so most models are able to identify the mistakes and their locations. However, they struggle to provide appropriate guidance without revealing the answer because the mistakes are generally related to quite basic operations like addition or multiplication, often in a one-step type of mathematical problems. Still, models like GPT-4 and Llama-3.1-405B are able to offer some reasonable guidance. In contrast, for MathDial, the contexts are longer, the mistakes are grounded in reasoning, and the responses are more structured. Yet, many LLMs do not meet the expectations for each dimension of the taxonomy, as discussed earlier. DAMR scores for Bridge and Mathdial are shown in Appendix Table 8 and 9, respectively. Combining both types of data in MRBench makes it both challenging and comprehensive.
@@
In summary, all LLMs and even human tutors lack some pedagogical abilities required for effective tutoring. While Llama-3.1-405B is the most effective, followed by Mistral and other state-of-the-art models, GPT-4 reveals the answer too quickly. Gemini is less coherent and accurate, and Sonnet focuses on human-likeness and encouraging tone but is less effective in other dimensions. Phi3 is the worst-performing model according to our analysis, as it fails to understand the context, while Llama-3.1-8B, despite being smaller, performs reasonably well. Human responses are also not perfect – Novice responses are ambiguous and short, whereas Expert responses are more focused on actionability and less on other dimensions. Overall, the proposed taxonomy precisely categorizes performance across 8 dimensions, reflecting the current state-of-the-art in AI tutors. Our study demonstrates that there is a considerable room for improvement in the pedagogical abilities of AI tutors.
@@
We also performed annotations using Prometheus2 and Llama-3.1-8B as critic LLMs. The correlation scores with human annotations are presented in Appendix Tables 5 and 6, respectively. Across both LLMs, it can be observed that most of the correlation scores are negative (except for the human-likeness dimension), indicating that the annotations from the LLMs are unreliable for the challenging pedagogical dimensions. There may be several reasons for this: (i) Prometheus2 is not trained on our taxonomy dimensions, except for the general human-likeness dimension, where the model shows slightly better correlations with positive scores. However, the score for human-likeness remains low and requires gold-standard responses, which are not unique and were unavailable in our case. (ii) We believe both LLMs have a limited understanding of rich pedagogical concepts, as they were not specifically trained on pedagogically rich datasets.
At the same time, we acknowledge that the experiments presented in this work are preliminary and have several limitations, including reliance on a specific prompt (see Figure 6) and the use of only two LLMs. Therefore, it is possible that better results may be achieved via extensive prompt engineering and experimentation with other LLMs as critics. We leave such experiments to future work.